### v1

# Reasoning on Time-Series for Financial Technical Analysis

Abstract.
While Large Language Models have been used to produce interpretable stock forecasts, they mainly focus on analyzing textual reports but not historical price data, also known as Technical Analysis. This task is challenging as it switches between domains: the stock price inputs and outputs lie in the time-series domain, while the reasoning step should be in natural language. In this work, we introduce Verbal Technical Analysis (VTA), a novel framework that combine verbal and latent reasoning to produce stock time-series forecasts that are both accurate and interpretable. To reason over time-series, we convert stock price data into textual annotations and optimize the reasoning trace using an inverse Mean Squared Error (MSE) reward objective. To produce time-series outputs from textual reasoning, we condition the outputs of a time-series backbone model on the reasoning-based attributes. Experiments on stock datasets across U.S., Chinese, and European markets show that VTA achieves state-of-the-art forecasting accuracy, while the reasoning traces also perform well on evaluation by industry experts.
Large Language Models, Time-Series, Technical Analysis
copyright: rightsretainedjournalyear: 2025conference: Workshop on Rethinking Financial Time-Series (ICAIF ’25); Nov 16 2025; Singapore, Singaporeccs: Computing methodologies Temporal reasoningccs: Applied computing Forecastingccs: Applied computing Economics
Table 1. Comparison of relevant works. Our work contributes a novel explainable financial signal for practitioners and produces some insights into how time-series forecasting can be made interpretable.
Models Domain Input Output
Financial LLMs
Fin-R1 (liu2025fin), Fino1 (qian2025fino1) Financial Question Answering Financial Reports Textual Answers
FinMem (yu2025finmem), FinCon (yu2024fincon) Investment Decision-Making Text, Tabular, Audio Binary (Buy/Sell)
GPT-4 (yu2023temporal), SEP (koa2024learning) Market Direction Forecasting News, Social Texts Binary or Quantile-Based
Time-Series LLMs
Time-LLM (jin2023time), CALF (liu2025calf) Time-Series Forecasting General Time-Series Time-Series Only
Ours
Verbal Technical Analysis (VTA) Financial Time-Series Forecasting Financial Time-Series Time-Series + Reasoning
1. Introduction
With the advent of Large Language Models (LLMs), an increasingly popular application is in financial analysis (wu2023bloomberggpt; xie2023pixiu). This spans a wide range of tasks, including financial question answering (FinQA) (liu2025fin; qian2025fino1), investment decision-making (yu2025finmem; yu2024fincon), and market forecasting (yu2023temporal; koa2024learning). Majority of existing approaches primarily utilize the strong natural-language capabilities of LLMs to analyze financial reports or do sentiment analysis on social texts (see Table 1), but neglects interpretable analysis on stock price data, which arguably contain useful information for financial practitioners. This is likely due to the fact that there are not many known techniques on time-series reasoning in the general domain (merrill2024language; chow2024towards). However, unlike other time-series data, financial time-series contains known interpretable signals which are widely studied by experts, known as Technical Analysis (kirkpatrick2010technical). We explore using these signals to verbally analyze financial time-series data, to produce stock forecasts with interpretable reasoning traces.
The use of LLMs for time-series reasoning is hindered by three main challenges. Firstly, current LLMs have limited capabilities in time-series forecasts. Some works have tackled this by modifying the embedding space to produce time-series outputs (jin2023time; liu2025calf), but this comes at the cost of interpretability as the LLM loses its natural language capability. Secondly, on a higher level, current LLMs are not known to have the ability to do verbal reasoning on time-series to produce accurate forecasts (merrill2024language; chow2024towards). This involves understanding how to best analyze the predictive signals in the time-series data in an unsupervised manner. Thirdly, the reasoning trace of the LLM further needs to be converted into time-series output to produce useful stock forecasts. LLMs are typically fine-tuned on next-token predictions (radford2019language), and using direct time-series outputs would not produce the best forecasts, which we verify empirically.
To address these problems, we present three key contributions. Firstly, we propose our Verbal Technical Analysis (VTA) framework, which combines a backbone time-series model (which we termed as “latent thinking”) with a reasoning LLM (termed as “verbal reasoning”) to produce interpretable stock time-series forecasts. This framework combines the strong pattern processing ability of state-of-the-art time-series models and the strong reasoning ability of LLMs to produce forecasts that are both accurate and interpretable. Secondly, for reasoning over time-series, the stock time-series data is converted into textual annotations (lin2024decoding) as inputs to the LLM. The reasoning trace is then optimized through a modified Group Relative Policy Optimization (GRPO) objective (shao2024deepseekmath) that uses an inverse Mean Squared Error (MSE) reward scoring, which we termed as Time-GRPO. Thirdly, to produce time-series outputs from the reasoning traces, we condition (ho2022classifier) the generated outputs from the time-series model on the reasoning-based attributes.
To demonstrate the effectiveness of VTA, we perform extensive experiments across established stock baselines (xu2018stock) and additional stock data across the U.S., Chinese, and European markets. We show that our model forecasts achieve state-of-the-art results in prediction accuracy, while also being interpretable. In addition, evaluation by industry experts show that the reasoning traces score high across various evaluation metrics from literature. Finally, to justify the practical capability of the model, we form Markowitz portfolios across the prediction length and show that the portfolios formed from VTA forecasts also perform well on investment metrics.
2. Related Works
Financial Large Language Models. The rise of Large Language Models (LLMs) has spurred a growing body of research on their application in finance. The earliest works focus on developing general-purpose financial LLMs, such as BloombergGPT (wu2023bloomberggpt) and FinMA (xie2023pixiu), by finetuning on a large set of financial corpora across multiple downstream tasks. Later works began to tackle the specific challenges of LLMs in finance. For example, works on financial question-answering (FinQA) (liu2025fin; qian2025fino1) focus on teaching LLMs to analyze financial reports, which requires the ability to read structured financial tables and extract insights from complex documents (zhu2021tat). LLMs for investment decision-making (yu2025finmem; yu2024fincon) typically utilize multi-agent systems to handle different parts of an investment decision, including but not limited to document analysis, memory, risk control, etc. LLMs for financial forecasting (yu2023temporal; koa2024learning) seeks to predict the direction in which the market will go. Typically, these works analyze textual sources in order to understand the sentiment or financial health of a company. Our work positions itself in this field by learning to produce interpretable forward-looking signals from financial time-series, which could benefit all applications above.
Time-Series Large Language Models. At the same time, there is also a growing body of research in utilizing LLMs in the time-series domain. Early works (xue2023promptcast; gruver2023large) utilized the use of their sequence completion capabilities by reframing the time-series forecasting task, prompting the LLM to complete a given string of numbers. Later works (jin2023time; liu2025calf) leveraged on their large scale parameters and robust pattern recognition by fine-tuning LLMs for time-series forecasting tasks. However, these approaches typically modify the embedding space of the LLMs, making them lose their original language reasoning capabilities. Some works have also started to explore the ability of LLMs to reason over time-series (kong2025time). It was found that language models are “remarkably bad” at zero-shot time-series reasoning (merrill2024language), whereas fine-tuning them using a latent encoder (chow2024towards) shows some promising early results in reasoning over time-series for captioning (but not forecasting). Our work builds on this line of research by exploiting the predictive signals in financial time-series to produce interpretable forecasts, providing some insights on the capabilities of LLMs in reasoning over time-series.
3. Verbal Technical Analysis
The Verbal Technical Analysis (VTA) framework is shown in Figure 1. There are three components: (1) In Time-Series Reasoning, we teach an LLM to verbally reason over the time-series inputs. This is done through a textual annotator to extract useful indicators, and a proposed Time-Series Group Relative Policy Optimization (Time-GRPO) method; (2) In Time-Series Forecasting, we train a backbone forecasting model, which can better learn from the complex low-level patterns in the time-series data; (3) In Joint Conditional Training, the time-series forecast is conditioned on the reasoning attributes, and the model is trained over the conditional and unconditional forecasts concurrently.
Figure 1. The Verbal Technical Analysis (VTA) framework. We first teach an LLM to reason over time-series data. The reasoning outputs are used to condition a time-series forecasting model, to produce forecasts with similar attributes. This results in forecasts with interpretable reasoning traces.
3.1. Problem Formulation
We consider the task of forecasting short-term stock prices, based on a historical window of $T$ days. Let $\mathbf{X}=\{\mathbf{x}_{t-T+1},\mathbf{x}_{t-T+2},\cdots,\mathbf{x}_{t}\}$, where the input vector consists of the open price, high price, low price, volume traded, closing price and adjusted closing price, i.e., $\mathbf{x}_{t}=\left[o_{t},h_{t},l_{t},v_{t},c_{t},p_{t}\right]$. We aim to generate an output $\mathbf{Y}=\{\mathbf{v},\mathbf{y}\}$, which consists of the verbal reasoning trace $\mathbf{v}$ and the price trajectory over the next $T^{\prime}$ trading days $\mathbf{y}=\{p_{t+1},p_{t+2},\cdots,p_{t+T^{\prime}}\}$.
3.2. Time-Series Reasoning
To teach an LLM to reason over time-series inputs, we use a textual annotator to extract useful interpretable signals for forecasting. The LLM uses these indicators to reason over the time-series to make forecasts without any supervision data. This is achieved through our proposed Time-Series Group Relative Policy Optimization (Time-GRPO) method, which uses a multi-stage reinforcement learning (RL) pipeline, together with a modified GRPO objective (shao2024deepseekmath).
The time-series input is first converted into textual annotations (lin2024decoding), which consist of its statistics information (jin2023time) (e.g., its mean, minimum and maximum values) and financial technical indicators (murphy1999technical) (e.g., moving averages, momentum, etc. ). Formally, we have:
(1) $\mathbf{X^{\prime}}=\mathbf{f}(\mathbf{X}),$
where $\mathbf{f}$ is the annotation function and $\mathbf{X^{\prime}}$ are the annotated values.
Training Objectives. Using both the time-series $\mathbf{X}$ and their annotations $\mathbf{X^{\prime}}$, we form a prompt $\mathbf{q}$, and let the LLM forecast the upcoming time-series sequence through verbal reasoning. The objective of the LLM is to produce an output $\mathbf{o}$, which consists of a sequence prediction $\mathbf{\hat{y}_{\theta}}$ and a verbal reasoning trace $\mathbf{\hat{v}_{\theta}}$. Formally, we denote the set of all task prompts as $\mathbf{\mathcal{Q}}$ and a group of generated outputs as $\mathbf{O}=\{\mathbf{o_{1}},\mathbf{o_{2}},\cdots,\mathbf{o_{G}}\}$. The time-series reasoning LLM policy $\pi_{\theta}$ is then optimized across all groups using the following Time-GRPO objective:
(2) $\mathcal{L}_{\text{time-grpo}}(\theta)=\mathbb{E}_{\mathbf{q}\sim\mathcal{Q},\{\mathbf{o_{i}}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\mathbf{O}|\mathbf{q})}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Big(\min\Big(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})}A_{i},\\ \text{clip}\Big(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})},1-\epsilon,1+\epsilon\Big)A_{i}\Big)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Big)\Bigg],\\ \!\!\!\!\!\!\!\!\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-\log\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-1,$
where $\epsilon$ and $\beta$ are hyper-parameters. $A_{i}$ denotes the advantage of the LLM policy, which is derived from a set of rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ that are associated with outputs $\mathbf{O}$ in each group:
(3) $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}.$
We utilize the format reward that was used in previous works (guo2025deepseek), that enforces the model to always employ a thinking process that is between <think><think> and </think></think> tags.
Ideally, the generated reasoning trace should also maximize the expected accuracy of the time-series forecasts. This is achieved by using the Mean-Squared Error (MSE) score as an additional reward:
(3) $r_{\text{MSE}}(\theta)=1/\left(\lambda\cdot\left\|\hat{\mathbf{y}}_{\theta}-\mathbf{y}\right\|_{2}^{2}\right),$
where $\lambda$ is a hyperparameter. The inverse MSE was used as the reward scores are to be maximized.
Training Pipeline. Following established practices in LLM fine-tuning literature (guo2025deepseek; ouyang2022training; lu2024deepseek; chow2024towards), Time-GRPO utilizes a multi-stage pipeline to fine-tune the time-series reasoning LLM.
The first stage represents the cold-start phase (guo2025deepseek). As there were no “gold-standard” supervision data, this stage is used for generating the initial training samples, guided by the $\mathcal{L}_{\text{time-grpo}}$ objective. Empirically, we find that the forecasting performance would not significantly improve here, but the process lets us generate training data for the next stage to fine-tune the base model.
The second stage focuses on teaching the model to produce more effective reasoning. This is achieved through rejection sampling, where we keep only the reasoning traces that lead to forecasts with lower Mean Squared Error (MSE). To ensure better consistency of training samples, we also bucket the samples across different stocks and time-periods, and filter for those with MSE in the bottom $10^{\text{th}}$ percentile in each bucket. The model is then trained on these filtered samples through the use of supervised fine-tuning (SFT).
The third stage optimizes the model for the best forecasting performance. Given that the model has learnt to reason over time-series data in the previous stage, this stage now aims to search for the best reasoning policy that can maximize the expected accuracy of the predicted time-series. For this stage, the model is also optimized using the $\mathcal{L}_{\text{time-grpo}}$ objective.
3.3. Time-Series Forecasting
To perform time-series forecasting, we employ LLM-based time-series models. Works have shown that the powerful contextual modeling capabilities of LLMs can be effectively adapted for time-series forecasting tasks (zhou2023one; chang2023llm4ts; cao2023tempo). A key technique is to align the time-series and language distributions (sun2023test; jin2023time) such that the model is able to understand the context of time-series data (e.g., up, down, steady, etc. ). For our backbone model, we repurpose an LLM for cross-modal fine-tuning (liu2025calf).
For this step, we first pass the time-series input $\mathbf{X}$ through an embedding layer, followed by a multi-head attention layer, to obtain the projected time tokens $\mathbf{X}_{\text{time}}$. Next, it is observed that similar words are usually close to each other in the LLM embedding space, and for non-text based tasks (e.g., time-series forecasting), it is sufficient to keep cluster centers to reduce training costs (sun2023test; liu2025calf). To do this, we perform Principal Component Analysis (PCA) to retrieve the principal word embeddings $\mathbf{\hat{D}}$. We then pass the projected time tokens $\mathbf{X}_{\text{time}}$ and the principal word embeddings $\mathbf{\hat{D}}$ through a Multi-head Cross-Attention layer. This lets us align the time tokens and word embeddings in the forecasting model’s embedding space to obtain the aligned cross-modal text tokens, i.e.,
(4) $\begin{gathered}\mathbf{X}_{\text{text}}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{C}}\right)\mathbf{V},\\ \text{where}\quad\mathbf{Q}=\mathbf{X}_{\text{time}}\mathbf{W}_{q},\quad\mathbf{K}=\hat{\mathbf{D}}\mathbf{W}_{k},\quad\mathbf{V}=\hat{\mathbf{D}}\mathbf{W}_{v}.\end{gathered}$
$\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ are the projection matrices for query, key and value in the multi-headed attention layer, and $C$ is the embedding dimension per attention head. $\mathbf{X}_{\text{text}}$ refers to the aligned text tokens.
Next, the projected time tokens $\mathbf{X}_{\text{time}}$ and aligned text tokens $\mathbf{X}_{\text{text}}$ are passed through consecutive LLM transformer blocks. To guide the modality alignment, after each transformer block in the temporal and textual branches, the embeddings are passed through a projection layer (chen2020simple) and matched via a feature regularization loss. This allows the forecasting model to tap into the deep contextual knowledge of the LLM while ensuring that the textual representations are consistent with the temporal dynamics at each layer. Formally, given $\mathbf{F}^{n}_{\text{time}}$ and $\mathbf{F}^{n}_{\text{text}}$, which are the outputs of the $n^{\text{th}}$ transformer block in the temporal and textual branches respectively, we define the feature regularization loss as:
(5) $\mathcal{L}_{\text{feature}}=\sum_{n=1}^{N}\gamma^{(N-n)}\,\text{sim}\bigl(\phi^{n}_{\text{text}}(\mathbf{F}^{n}_{\text{text}}),\phi^{n}_{\text{time}}(\mathbf{F}^{n}_{\text{time}})\bigr),$
where $\gamma^{(N-n)}$ are the scaling hyperparameters, $\text{sim}(\cdot,\cdot)$ is the $L_{1}$ regularization loss to ensure embedding similarity, and $\phi^{n}_{\text{time}}$, $\phi^{n}_{\text{text}}$ are the projection layers in the temporal and textual branches.
At the end of the transformer blocks, the features are passed through a final dense layer to produce the temporal- and text-based outputs, $\hat{\mathbf{y}}_{\text{time}}$ and $\hat{\mathbf{y}}_{\text{text}}$. These are also matched via $L_{1}$ loss:
(6) $\mathcal{L}_{\text{output}}=\text{sim}(\hat{\mathbf{y}}_{\text{time}},\hat{\mathbf{y}}_{\text{text}}).$
The temporal-based output $\hat{\mathbf{y}}_{\text{time}}$ is extracted, which we denote as the time-series forecast $\hat{\mathbf{y}}_{\phi}(\mathbf{X})$.
3.4. Joint Conditional Training
On its own, the time-series forecasting pipeline represents a black-box model, given that the embedding space of the LLM blocks have been modified, resulting in only time-series outputs. To preserve the interpretability of the time-series forecasts, we condition the time-series forecasts on the outputs produced by the reasoning model. At the same time, to also preserve the forecast accuracy, we fine-tune the model to optimize for both the conditional and unconditional forecasts concurrently.
Figure 2. Joint conditional training component.
For this step, we first prompt for the reasoning output $\mathbf{o}$ using the previously fine-tuned time-series reasoning policy $\pi_{\theta}$. Next, we extract descriptive attribute classes $\mathbf{c}$ (i.e., its maximum, minimum, mean values) from the generated time-series $\hat{\mathbf{y}}_{\theta}$, which are analogous to the conditional class labels in text-conditional image generation (dhariwal2021diffusion). These are then used to condition the time-series forecasts via joint conditional training (see Figure 2): For each label, we concatenate it with the time-series forecasts $\hat{\mathbf{y}}_{\phi}(\mathbf{X})$ from the time-series forecasting model. These individual inputs pass through separate linear layers for fine-tuning, and are then aggregated via a projection layer to generate the conditioned time-series forecasts, which we denote here as $\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})$.
Finally, we use a single neural network to parameterize both the conditional and unconditional models (ho2022classifier). Both the unconditional and conditional pipelines are trained concurrently by randomly setting $\mathbf{c}$ to the unconditional class identifier $\mathbf{\varnothing}$ with a predefined probability $p_{\text{uncond}}$, set as a hyperparameter. The overall model is then fine-tuned using MSE loss with the ground truth series $\mathbf{y}$:
(7) $\mathcal{L}_{\text{forecast}}(\phi)=\mathbb{E}_{\mathbf{X},\mathbf{y},\mathbf{c}}\left[\left\|\hat{\mathbf{y}}_{\psi}(\mathbf{X},\tilde{\mathbf{c}})-\mathbf{y}\right\|^{2}\right],$
(8) $\tilde{\mathbf{c}}\sim\begin{cases}\mathbf{c},&\text{with probability }1-p_{\text{uncond}}\\ \mathbf{\varnothing},&\text{with probability }p_{\text{uncond}}\end{cases}.$
During inference, our forecast is then a combination of the conditional and unconditional forecasts:
(9) $\hat{\mathbf{y}}=\hat{\mathbf{y}}_{\phi}(\mathbf{X})+s\cdot\bigl(\hat{\mathbf{y}}_{\psi}(\mathbf{X},\mathbf{c})-\hat{\mathbf{y}}_{\phi}(\mathbf{X})\bigr),$
where $s$ is a hyperparameter representing the guidance scale, that controls the reasoning guidance.
4. Experiments
Dataset: We evaluate our Verbal Technical Analysis (VTA) model extensively across multiple datasets. The first is the ACL18 StockNet dataset (xu2018stock), which includes historical price data for 88 U.S. stocks that are selected to represent the top 8-10 companies by market capitalization in each of the major industries. The dataset spans the period of 01/09/2012 to 01/09/2017. This dataset is a standard stock prediction benchmark that has been evaluated in multiple works (koa2023diffusion; feng2018enhancing; sawhney2020deep; feng2021time; li2023pen; chen2025dhmoe).
To further show the generalization ability of the model, we also collect additional stock data from across the U.S., Chinese, and European markets for testing. To ensure a bias-free selection, we choose the stocks from well-known indices, i.e., the Dow Jones, the FTSE China A50 Index and the EURO STOXX 50. For these datasets, we test on the time period from 01/01/2024 to 01/01/2025.
Baselines: We compare against 12 state-of-the-art time-series methods: Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Autoformer (wu2021autoformer), DLinear (zeng2023transformers), FiLM (zhou2022film), Crossformer (zhang2023crossformer), MICN (wang2023micn), LightTS (campos2023lightts), TimesNet (wu2022timesnet), TSMixer (chen2023tsmixer) and Non-Stationary Transformer (liu2022non). We also compare with two LLM-based time-series models: TimeLLM (jin2023time) and CALF (liu2025calf). These models are not explainable, including the two LLM-based models, which modify the embedding space for forecasting.
For evaluation against explainable models, we compare with reasoning LLMs: GPT-4.1 mini (openai2025gpt41) and DeepSeek-R1 (guo2025deepseek). To do so, we prompt these models to produce the time-series forecasts (gruver2023large; wang2024news) by reasoning on the time-series inputs.
Table 2. Performance comparison. The best baselines are underlined, and the best results are bolded.
StockNet China A50 EUROSTOXX 50 Dow Jones All Data % Improvement
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Large Language Models
GPT-4.1 mini (openai2025gpt41) 0.0846 0.1827 0.4875 0.3191 0.0997 0.2128 0.1340 0.2358 0.2014 0.2376 0.4153 0.1072
DeepSeek-R1 (guo2025deepseek) 0.0788 0.1853 0.2920 0.3093 0.0776 0.2095 0.1227 0.2251 0.1428 0.2323 0.1750 0.0868
Time-Series Models
Informer (zhou2021informer) 2.1846 0.9778 4.6823 1.1080 2.8986 1.0968 3.7031 1.1255 3.3672 1.0770 0.9650 0.8030
Transformer (vaswani2017attention) 1.5071 0.7762 3.9865 0.9784 2.1002 0.9042 2.8248 0.9194 2.6047 0.8946 0.9548 0.7628
Crossformer (zhang2023crossformer) 1.1848 0.6475 4.3396 0.9641 1.5656 0.7644 2.3503 0.8173 2.3601 0.7983 0.9501 0.7343
TSMixer (chen2023tsmixer) 1.4974 0.8193 3.5863 1.0905 1.1696 0.7227 2.5077 0.8473 2.1902 0.8700 0.9462 0.7561
Reformer (kitaev2020reformer) 1.1823 0.7628 2.4537 0.8503 1.5342 0.8302 2.4280 0.9048 1.8995 0.8370 0.9380 0.7465
LightTS (campos2023lightts) 0.6081 0.5213 1.0634 0.5509 0.6160 0.5128 1.0994 0.6103 0.8467 0.5488 0.8609 0.6134
DLinear (zeng2023transformers) 0.1589 0.3021 0.3880 0.4126 0.1716 0.3189 0.2407 0.3566 0.2398 0.3475 0.5088 0.3896
FiLM (zhou2022film) 0.0806 0.1927 0.2852 0.3085 0.0894 0.2157 0.1246 0.2370 0.1449 0.2385 0.1873 0.1103
Non-stationary (liu2022non) 0.0729 0.1822 0.2723 0.2993 0.0861 0.2079 0.1207 0.2305 0.1380 0.2300 0.1463 0.0776
MICN (wang2023micn) 0.0764 0.1878 0.2498 0.2922 0.0874 0.2108 0.1373 0.2405 0.1377 0.2328 0.1449 0.0888
Autoformer (wu2021autoformer) 0.0748 0.1866 0.2427 0.2947 0.0853 0.2103 0.1132 0.2273 0.1290 0.2297 0.0868 0.0765
TimesNet (wu2022timesnet) 0.0708 0.1789 0.2527 0.2902 0.0796 0.2022 0.1112 0.2203 0.1286 0.2229 0.0841 0.0482
Time-Series LLMs
TimeLLM (jin2023time) 0.0704 0.1780 0.2439 0.2850 0.0776 0.1993 0.1127 0.2216 0.1262 0.2210 0.0663 0.0400
CALF (liu2025calf) 0.0674 0.1738 0.2412 0.2843 0.0762 0.1957 0.1092 0.2181 0.1235 0.2180 0.0460 0.0267
Our Model Avg Avg
VTA (Ours) 0.0659 0.1701 0.2265 0.2737 0.0748 0.1929 0.1040 0.2120 0.1178 0.2122 0.4672 0.3417
Implementation Details: All LLMs used in VTA, including the reasoning model and transformer blocks for time-series forecasting, are trained using Low-Rank Adaptation (LoRA) (hu2022lora). Both input and output lengths $T$ and $T^{\prime}$ are set to 10, which is considered short-term forecasting in time-series works (li2022generative; liu2025calf). Technical Analysis is typically utilized for short-term stock trading (schwager1995technical).
For the reasoning model, we use Qwen2.5-7B-Instruct (qwen2.5) as our base model. For the forecasting model, we use GPT-2 (radford2019language) as the base model. For hyperparameters, we set the conditional probability $p_{\text{uncond}}$ to 0.3 and the guidance scale $s$ to 0.1.
5. Results
Performance Comparison. Table 2 reports the forecasting results.
  • The inference-only reasoning LLMs (i.e., GPT-4.1 mini, DeepSeek-R1) do not show very strong performances, as they are likely not fine-tuned for time-series forecasting. However, they were still able to beat some of the fine-tuned time-series models (e.g., Transformer, DLinear), which demonstrate some effectiveness of verbally reasoning over time-series inputs to do forecasting.
  • Among the time-series baselines, the biggest performance jump comes from the models which decomposes the data by trend and seasonality (i.e., FiLM, MICN, Autoformer and TimesNet). This could be attributed to the characteristics of stock prices, which are often affected by long-term and short-term business cycles (dalio2018principles). Another notable model that performs well is the non-stationary transformer, which might be attributed to the non-stationary behavior of stock price data (malkiel2019random).
  • The time-series LLMs are able to surpass the performance of non-LLM-based models. These models typically align the LLM’s internal word embeddings to time-series embeddings to do time-series forecasting. It is possible that the intrinsic knowledge of the LLM helps the model to understand the characteristics of stock data, thus allowing it to capture forecasting patterns better.
  • Our proposed VTA model demonstrates the best performance in stock forecasting in both MSE and MAE. VTA explicitly combines internal (latent) understanding with external (verbalized) reasoning. The empirical improvements suggest that integrating these two techniques can be beneficial for time-series forecasting. In addition to improved accuracy, the VTA model also produces interpretable reasoning traces for its forecasts, which do not exist in most baseline models.
Figure 3. Correctness reward over steps.
Table 3. Ablation study of the LLM fine-tuning stages.
Llama-3.1-8B Qwen-2.5-3B Qwen-2.5-7B (Ours)
Base Model MSE 0.1482 0.1707 0.0949
MAE 0.2543 0.2181 0.2040
Cold Start RL MSE 0.1475 0.1648 0.0941
MAE 0.2536 0.2153 0.2036
SFT for Reasoning MSE 0.1168 0.1032 0.0893
MAE 0.2267 0.1884 0.1997
RL for Reasoning MSE 0.0955 0.0832 0.0686
MAE 0.2062 0.1745 0.1741
Conditioning (VTA) MSE 0.0667 0.0672 0.0659
MAE 0.1713 0.1710 0.1701
Figure 4. An example of the reasoning trace and forecasts produced by our VTA framework.
Ablation Study. We conduct an ablation study to demonstrate the effectiveness of the model design.
  • From Figure 3, we observe that the inverse MSE reward $r_{\text{MSE}}$, a component of the Time-GRPO objective, increases across the number of training steps. This suggests that it is possible to learn verbal reasoning steps for time-series forecasting, which our VTA model was able to achieve.
  • From Table 3, we see that each fine-tuning stage helps to improve the results of the model. However, the first RL fine-tuning, which uses the Time-GRPO objective, was not efficient by itself, showing a small improvement of 1.6% over the base model in MSE averaged across all variants.
  • However, after rejection sampling and SFT to teach the model how to reason over time-series, the second RL fine-tuning, which uses the same Time-GRPO objective, produces an average improvement of 20.3%. This highlights the usefulness of fine-tuning over a multi-stage pipeline.
  • Finally, conditioning on the additional forecasting model helped to improve the performance further, showing the effectiveness of our combined verbal and latent VTA framework.
Reasoning Quality. To evaluate the quality of the reasoning traces, we refer to past works on LLM explainability (koa2024learning; lin2024decoding) to design a set of relevant metrics for our task. Using these metrics, we surveyed 25 industry experts with professional experience in financial market analysis, with backgrounds from organizations such as JPMorgan, UBS, Evercore, and Allianz Global Investors, etc. For each, we presented the model outputs from VTA (ours), GPT-4.1 mini, and Deepseek-R1, showing both the forecasts and the textual reasoning. Respondents were blind to which model produced which output and were shown 15 randomly selected samples for evaluation. The samples were rated from 1 to 5.
Table 4. Comparison on reasoning quality.
Clarity Depth Accuracy Coherence Relevance
GPT-4.1 mini 3.43 1.88 2.15 3.15 2.53
Deepseek R1 3.07 2.20 2.36 3.03 2.56
VTA (Ours) 3.74 4.04 3.37 3.60 3.71
From Figure 4, the results show that VTA achieved the highest average rating across all five metrics when compared to the other two LLMs. The best performance gains were observed in Depth, Accuracy, and Relevance. These criteria most directly reflect technical reasoning ability and the use of financial indicators. These findings suggest that our model, which was designed specifically for technical analysis, was able to produce reasoning outputs that were preferred by domain experts, demonstrating its practical strengths. The differences in Coherence and Clarity were smaller, which can be attributed to the fact that general-purpose LLMs like GPT-4.1 mini and Deepseek-R1 are good at producing fluent, well-structured text, even if they lack domain specialization. An example of the VTA reasoning outputs can be found in Figure 4.
Portfolio Optimization. To justify the practical capability of VTA, we also evaluate the model in a real-life investment setting. Given that the model produces multi-step predictions, we are able to form portfolios that maximize the returns and minimize the volatility across the prediction length (koa2023diffusion). This is achieved by performing Markowitz optimization (selection1952harry) across the 10-day predictions. The portfolio is rebalanced daily, using the predicted returns and their covariance matrix.
For evaluation, we compare against similar portfolios formed using the top-5 performing time-series models and all LLM baselines. The portfolio are compared on common investment metrics, such as their returns, volatility, maximum drawdown and Sharpe ratio (sharpe1994sharpe), which is a measure of risk-adjusted returns. We evaluate the portfolios across all 4 datasets, and report the average metrics.
From Table 5, we see that the portfolio constructed using VTA predictions demonstrates strong overall performance. It ranks as the strongest baseline on returns, volatility and maximum drawdown, and the values are very close to the top-performing models for each. Notably, VTA achieves the highest Sharpe ratio among all models, which represents the risk-adjusted returns. Given that the Sharpe ratio is one of the most common measure of investment performance, this highlights the practical effectiveness of VTA.
Table 5. Comparison on common portfolio metrics.
Returns Volatility Drawdown Sharpe
GPT-4.1 mini 0.1868 0.1226 -0.0947 1.3096
Deepseek-R1 0.2069 0.1356 -0.1243 1.4074
FiLM 0.2211 0.1184 -0.1085 1.4421
Non-stationary 0.2122 0.1186 -0.0825 1.4430
MICN 0.1603 0.1193 -0.1094 1.1809
Autoformer 0.2495 0.1341 -0.1121 1.4736
TimesNet 0.1714 0.1198 -0.0947 1.2748
TimeLLM 0.2185 0.1193 -0.1040 1.5230
CALF 0.2019 0.1247 -0.0981 1.4566
VTA (Ours) 0.2409 0.1185 -0.0883 1.7190
6. Conclusion
In this work, we tackled the task of verbally reasoning over financial time-series data. This task is challenging as it switches between the time-series and natural language domain for the stock price data and the reasoning step. To deal with this, we introduce our Verbal Technical Analysis (VTA) framework, which combines verbal and latent reasoning to produce interpretable time-series forecasts. The framework utilizes our Time-GRPO method to finetune the reasoning model, and conditions its forecasts on the reasoning attributes. We conducted extensive experiments across multiple benchmarks and find VTA can achieve state-of-the-art forecasting accuracy while also producing high-quality reasoning traces.
Appendix A List of Technical Indicators
Table 6. A list of the financial technical indicators used in Time-GRPO. Indicator Description Formula Simple Moving Average (SMA) Identifies trend by smoothing price data over a period $\text{SMA}=\frac{1}{n}\sum_{i=1}^{n}\text{Price}_{i}$ Exponential Moving Average (EMA) Measures the trend by smoothing price data with greater weight to recent prices $\text{EMA}_{t}=\text{Price}_{t}\cdot\alpha+\text{EMA}_{t-1}\cdot(1-\alpha)$ Momentum Tracks the speed of price changes for trend momentum $\text{Momentum}=\text{Close}_{t}-\text{Close}_{t-n}$ Relative Strength Index (RSI) Identifies overbought/oversold conditions using average gains and losses $\text{RSI}=100-\left(\frac{100}{1+\frac{\text{Avg Gain}}{\text{Avg Loss}}}\right)$ MACD Line Measures momentum via difference between short and long EMAs $\text{MACD}=\text{EMA}_{12}-\text{EMA}_{26}$ Williams %R Measures overbought/oversold conditions by comparing close to recent highs $\text{Williams}~\%R=\frac{\text{Highest High}_{n}-\text{Close}_{t}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times(-100)$ Commodity Channel Index (CCI) Identifies price deviations from a moving average for cyclical trends $\text{CCI}=\frac{\text{Price}-\text{MA}}{0.015\times\text{Mean Deviation}}$ Average Directional Index (ADX) Measures trend strength using directional movement $\text{ADX}=100\times\frac{\text{EMA}(|\text{DM}^{+}-\text{DM}^{-}|)}{\text{DM}^{+}+\text{DM}^{-}}$ Bollinger Bands Measures volatility using standard deviations around a moving average $\text{Upper Band}=\text{MA}+k\cdot\sigma$ $\text{Lower Band}=\text{MA}-k\cdot\sigma$ Stochastic Oscillator Measures momentum by comparing current close to a range of highs and lows $\%K=\frac{\text{Close}_{t}-\text{Lowest Low}_{n}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times 100$
Appendix B Evaluation Metrics
We evaluated the quality of the reasoning traces produced by our model using five criteria: clarity, depth, accuracy, coherence, and relevance. Each reasoning trace and its associated numeric forecast was scored on a scale from 1 (poor) to 5 (excellent). The criteria are defined as follows:
  • Clarity: How clearly and succinctly the reasoning explains its analysis in a structured manner.
  • Depth: How well the reasoning incorporates explicit financial or technical indicators (e.g., MACD, RSI, Bollinger Bands, EMA) to meaningfully support its conclusions.
  • Accuracy: How precisely financial indicators are interpreted and technically described.
  • Coherence: How logically consistent and organized the reasoning is, ensuring clear alignment between analysis and conclusion.
  • Relevance: How directly and effectively the chosen indicators are linked to the stock-price forecast provided.